OpenStudy (anonymous):

-I3y+5I+6=2 can anybody help me get started??

OpenStudy (anonymous):

No problem. For any linear equation, you must try to isolate y. So, subtract 6 from both sides. Then, divide by negative one on both sides to get rid of the negative symbol from outside the absolute function. When you hit the absolute function, you must separate the two equations, like this: 3y +5 = -2 AND 3y +5 = 2 And solve like a normal equation from there, subtracting the constants, and solving for x. Does that help?

OpenStudy (anonymous):

if you subtract 6 from both sides you get\[-|3y+5|=-4\] or \[|3y+5|=4\]

OpenStudy (anonymous):

That's right, forgot to subtract 6 from both sides in my first step. Thanks, satellite.

OpenStudy (anonymous):

@nickjesus3 thank you so much it sure does..........this is almost like spanish i havent seen this in years

OpenStudy (anonymous):

then you can solve \[3y+5=4\] or \[3y+5=-4\] to get your answers

OpenStudy (anonymous):

thanks to the both of you

OpenStudy (anonymous):

im not seeing how we get 4 or -4

OpenStudy (anonymous):

@nickjesus3 can you explain that to me

OpenStudy (anonymous):

Yeah, so when you started off with -I3y+5I+6=2, you needed to subtract 6 from the right and left side of the equation. This is done so that, eventually, the y-term will be by itself (as in y=5 or y=4, whatever it is)

OpenStudy (anonymous):

would it look like that

OpenStudy (anonymous):

It would have 4 and negative 4, yes. The reason why the 4 on can be negative or positive is because of the absolute value function.

OpenStudy (anonymous):

This guy explains it here: http://patrickjmt.com/solving-absolute-value-equations-ex-1/

OpenStudy (anonymous):

thanks